OpenStudy (jkbo):

please help. find the 23rd term in the following sequence 11,14,17,20,23

OpenStudy (anonymous):

11 + (22)3 = 11 + 66 = 77

OpenStudy (anonymous):

I used the formula a_n = a + (n-1)d where d = 3, n = 23

OpenStudy (anonymous):

a = 11

OpenStudy (jkbo):

thank you so much